OpenStudy (mony01):

By using integration by parts. Solve the solution for the following DE. y'=(cos^(2)y)ln(t)

OpenStudy (tkhunny):

Are you sure it is "Integration by Parts" and not "Separation of Variables"? \(\dfrac{dy}{\cos^{2}(y)} = ln(t)\;dt\) \(\sec^{2}(y)\;dy = ln(t)\;dt\) Well, okay, the right hand side will still take Integration by parts.

OpenStudy (mony01):

is the right hand side this \[tln(t)-\int\limits(t)\frac{ 1 }{ t}dt\]

OpenStudy (tkhunny):

Looks good to me. Why do you doubt?

OpenStudy (tkhunny):

What did you get for the left side?

OpenStudy (mony01):

tan(y)=tln(t)-t+c